# Gautschi-type and implicit–explicit integrators for constrained wave equations

### v1

@@ 2.2. Existence of solutions
Before introducing numerical schemes for constrained systems of wave type, we need to discuss the existence of solutions to (2.1). Throughout this section, we denote the well-known Sobolev-Bochner spaces for a Banach space $X$ by $L^{2}(0,T;X)$ and $H^{1}(0,T;X)$; cf. [Zei90, Ch. 23]. Moreover, $C([0,T],X)$ denotes the space of continuous functions in $[0,T]$ with values in $X$. As a first step, we give a classical existence result for the linear case without constraints. Here, linear means that the right-hand side $f$ is independent of $u$.
@@
with exponent $\kappa\coloneqq C_{2}/\sqrt{\mu_{1}}$ and the problem-dependent norm $\|\bullet\|^{2}_{\mathcal{A}_{1}}\coloneqq\langle\mathcal{A}_{1}\,\bullet\,,\bullet\rangle$.
@@
By an adaption of [Zei90, Cor. 24.3], the problem has a unique solution $u$ within the given spaces. It remains to show the stability estimate (2.6). For this, note that $u$ satisfies the system
@@
with the same initial values and right-hand side $\tilde{f}=f-\mathcal{A}_{2}u\in L^{2}(0,T;\mathcal{H}^{*})$. Along the same lines as in [LM72, Ch. 3, Lem 8.3] and Young’s inequality, we conclude that
@@
Consider Assumptions 2.1, 2.2, 2.3, and 2.6 as well as a right-hand side $g\in H^{2}(0,T;\mathcal{Q}^{*})$. Further, assume consistent initial data
@@
Then there exists a $T>0$ and a unique solution to (2.1) on $[0,T]$ with $u(0)=u^{0}$, $\dot{u}(0)=w^{0}$ satisfying
@@
Again, the decomposition $u=u_{\text{ker}}+u_{\text{c}}$ yields that we only have to study the part in the kernel and we proceed as in Lemma 2.10. We employ estimate (2.7) to set up a fixed-point argument with the nonlinearity
@@ 2.3. Spatial discretization
where the right-hand sides are now functions $f\colon[0,T]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g\colon[0,T]\to\mathbb{R}^{m}$. Moreover, we assume consistent initial values, i.e., $x(0)=x^{0}$, $\dot{x}(0)=y^{0}$ with $Bx^{0}=g(0)$ and $By^{0}=\dot{g}(0)$.
Remark 2.12.
Due to the semi-explicit structure of semi-discrete system, one can easily see that the stated assumptions imply that the DAE (2.8) is of index $3$; cf. [HW96, Ch. VII.1].
@@ 3. An implicit–explicit Crank–Nicolson Scheme
This section is devoted to the extension of the IMEX scheme introduced in [HL21] to constrained systems. For this, we consider a uniform partition of the time interval $[0,T]$ into $0=t_{0}<t_{1}<\dots<t_{N}=T$ with constant time step size $\tau$, i.e., $t_{n}=n\,\tau$. Furthermore, we denote by $u^{n}$ the computed approximation of $u$ at time $t_{n}$ and abbreviate function evaluations by $g^{n}\coloneqq g(t_{n})$. Throughout, we use the notation
@@ 3.1. Review of the unconstrained case
with initial conditions $u(0)=u^{0}$, $\dot{u}(0)=w^{0}$. A Crank–Nicolson discretization, where the nonlinear term is treated by the left and right rectangle rule (see [HL21] for more details) yields the IMEX scheme
@@
Theorem 3.1 (Second-order convergence, unconstrained case [HL21, Th. 2.9]).
@@
Let us already mention that in the case of $\mathcal{D}=0$, one can derive an equivalent two-step formulation, which eliminates $w^{n}$, see Lemma 3.3 below.
@@ 3.2. An implicit–explicit scheme of second order
Second, we obtain $u^{n+1}\in\mathcal{V}$ simply by
@@
Here, the existence of a unique solution $w^{n+1}\in\mathcal{H}_{\ker}+\mathcal{V}_{\text{c}}$ is only guaranteed if $\mathcal{H}_{\ker}$ is a proper subspace of $\mathcal{H}$; cf. [Zim21, Lem. 9.22]. In this case, the expression $\mathcal{B}w^{n+1}$ is indeed well-defined, even though $w^{n+1}\not\in\mathcal{V}$. In the following, we assume the unique solvability of (3.2d)-(3.2e) such that the IMEX scheme (3.2) is well-defined overall.
@@
$\quad=-\frac{\tau^{2}}{4}\mathcal{A}\,\big{(}\tau w^{n+1/2}+\tau w^{n-1/2}\big{)}-\frac{\tau}{2}\,\mathcal{A}\big{(}u^{n}+u^{n-1}\big{)}+\tau f^{n}-\tau\mathcal{B}^{*}\big{(}\tfrac{\tau}{2}\,\lambda^{n+1}+\tfrac{\tau}{2}\,\lambda^{n}+\mu^{n}\big{)}$
$\quad=-\frac{\tau^{2}}{4}\mathcal{A}\,\big{(}u^{n+1}+2u^{n}+u^{n-1}\big{)}+\tau f^{n}-\tau\mathcal{B}^{*}\big{(}\tfrac{\tau}{2}\,\lambda^{n+1}+\tfrac{\tau}{2}\,\lambda^{n}+\mu^{n}\big{)}.$
@@ 3.3. Convergence analysis
Similarly, we have by (3.2e),
@@
$=w_{\text{ker}}(t_{n})-w_{\text{ker}}^{n}+\mathcal{B}^{-}\bigl{(}g^{n}-\tfrac{1}{\tau}\,(g^{n+1/2}-g^{n-1/2})\bigr{)}$
@@
We aim to show a similar error representation for the parts in the kernel of $\mathcal{B}$ as in [HL21]. For technical reasons, we introduce the following projection operator.
@@
With $\mathcal{D}_{\ker}$ from Section 2.1, we introduce the operators $\mathcal{R}_{\pm}$ and $\mathcal{R}$ by
The essential properties of these operators are collected in the following lemma.
Lemma 3.7.
$\tau\mathcal{R}_{+}^{-1}\begin{bmatrix}-x\\ \mathcal{D}_{\ker}x\end{bmatrix}=(\operatorname{id}-\mathcal{R})\begin{bmatrix}0\\ x\end{bmatrix}$
and, thus, the bound
Proof.
The bounds on $\mathcal{R},\mathcal{R}_{+}^{-1}$ follow from [HL21, Lem. 2.4]. From the definition of $\mathcal{R}_{\pm}$, we have
and multiplying with $\mathcal{R}_{+}^{-1}$ gives the claim. ∎
@@
where we have the two defects $\delta^{n}_{u}$ and $\delta^{n}_{w}$, which essentially come from the trapezoidal rule and satisfy $\|\delta^{n}_{u}\|_{\mathcal{V}_{\ker}}+\|\delta^{n}_{w}\|_{\mathcal{H}_{\ker}}\leq C\tau^{3}$. For $\widehat{f}^{n}=f(t_{n},u(t_{n}))$ and $f^{n}=f(t_{n},u^{n})$, the other terms are given by
@@
Resolving the error recursion in Lemma 3.8, and using the fact that $(\operatorname{id}-\mathcal{R})$ allows for summation by parts, the proofs of Theorems 2.9 and 3.3 in [HL21] also apply to the present case. ∎
@@
We introduce the notation $\widehat{u}_{\text{ker}}^{n}=u_{\text{ker}}(t_{n})$, $\widehat{u}_{\text{c}}^{n}=u_{\text{c}}(t_{n})$ and further denote with $\delta^{n}_{u}$, $\delta^{n}_{w,1}$ the defects from the trapezoidal rule which satisfy $\|\delta^{n}_{u}\|_{\mathcal{V}}+\|\pi_{\text{ker}}\delta^{n}_{w,1}\|_{\mathcal{H}}\leq C\tau^{3}$. We can thus conclude that the exact solution satisfies
@@ 4. A Gautschi-type Integrator
This section is devoted to the extension of Gautschi-type integrators to constrained systems. For this, we restrict ourselves to the damping-free free case, i.e., we set $\mathcal{D}=0$. Moreover, we consider the stronger assumption on the nonlinearity, i.e., Assumption 2.6 (a).
@@ 4.1. Review of the unconstrained case
We recall solution strategies for conservative second-order systems as discussed in [HL99, Gri02]. Considering the corresponding first-order formulation, we consider the equations
@@
With the solution formula (4.4), we conclude for the solution of (4.1) with a constant right-hand side, i.e., for $f(t,u)\equiv f$, that
$u(t+\tau)+u(t-\tau)$ $=2\,\cos(\tau\Omega)\,u(t)+2\int_{0}^{\tau}(\tau-s)\,\operatorname{sinc}((\tau-s)\,\Omega)\,f\,\mathrm{d}s$
$=2\,\cos(\tau\Omega)\,u(t)+2\,\mathcal{A}^{-1}\big{[}1-\cos(\tau\Omega)\big{]}\,f$
@@
Following [HL99], this may also be written in the alternative form
@@ 4.2. An implicit solution formula
We come back to the constrained case and consider the PDAE (2.1). In the linear case, i.e., if $f(t,u)\equiv f(t)$, we can explicitly write down the solution using the decomposition of $\mathcal{V}$ from (2.3), the first-order formulation, and the variation-of-constants formula; cf. [EM13]. This solution formula is also applicable to the general semi-linear case, leading to an implicit formula. As in the unconstrained case, quadrature rules then yield numerical schemes.

### v2

@@ 2.2. Existence of solutions
Before introducing numerical schemes for constrained systems of wave type, we need to discuss the existence of solutions to (2.1). Throughout this section, we denote the well-known Sobolev-Bochner spaces for a Banach space $X$ by $L^{2}(0,T;X)$ and $H^{1}(0,T;X)$; cf. [Zei90a, Ch. 23]. Moreover, $C([0,T],X)$ denotes the space of continuous functions in $[0,T]$ with values in $X$. As a first step, we give a classical existence result for the linear case without constraints. Here, linear means that the right-hand side $f$ is independent of $u$.
@@
with exponent $\kappa\coloneqq C_{2}/\sqrt{\mu_{1}}$ and the problem-dependent norm $\|\bullet\|^{2}_{\mathcal{A}_{1}}\coloneqq\langle\mathcal{A}_{1}\,\bullet\,$$\bullet\rangle$.
@@
By an adaption of [Zei90a, Cor. 24.3], the problem has a unique solution $u$ within the given spaces. It remains to show the stability estimate (2.6). For this, note that $u$ satisfies the system
@@
with the same initial values and right-hand side $\tilde{f}=f-\mathcal{A}_{2}u\in L^{2}(0,T;\mathcal{H}^{*})$. Along the same lines as in [LioM72a, Ch. 3, Lem 8.3] and Young’s inequality, we conclude that
@@
Consider Assumptions 2.1, 2.2, 2.3, and 2.6 as well as a right-hand side $g\in H^{2}(0,\widetilde{T};\mathcal{Q}^{*})$. Further, assume consistent initial data
@@
Then there exists a $0<T\leq\widetilde{T}$ and a unique solution to (2.1) on $[0,T]$ with $u(0)=u^{0}$, $\dot{u}(0)=w^{0}$ satisfying
@@
Again, the decomposition $u=u_{\text{ker}}+u_{\text{c}}$ yields that we only have to study the part in the kernel, and we proceed as in Lemma 2.11. We employ estimate (2.7) to set up a fixed-point argument with the nonlinearity
@@
where the right-hand sides are now functions $f\colon[0,T]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g\colon[0,T]\to\mathbb{R}^{m}$. Moreover, we assume consistent initial values, i.e., $x(0)=x^{0}$, $\dot{x}(0)=y^{0}$ with $Bx^{0}=g(0)$ and $By^{0}=\dot{g}(0)$. Due to the semi-explicit structure of semi-discrete system, one can easily see that the stated assumptions imply that the DAE (2.8) is of index $3$; cf. [HaiW96, Ch. VII.1].
@@ 3. An implicit–explicit Crank–Nicolson Scheme
This section is devoted to the extension of the IMEX scheme introduced in [HocL21] to constrained systems. For this, we consider a uniform partition of the time interval $[0,T]$ into $0=t_{0}<t_{1}<\dots<t_{N}=T$ with constant time step size $\tau$, i.e., $t_{n}=n\,\tau$. Furthermore, we denote by $u^{n}$ the computed approximation of $u$ at time $t_{n}$ and abbreviate function evaluations by $g^{n}\coloneqq g(t_{n})$. Throughout, we use the notation
@@ 3.1. Review of the unconstrained case
with initial conditions $u(0)=u^{0}$, $\dot{u}(0)=w^{0}$. A Crank–Nicolson discretization, where the nonlinear term is treated by the left and right rectangle rule (see [HocL21] for more details) yields the IMEX scheme
@@
Theorem 3.1 (Second-order convergence, unconstrained case [HocL21, Th. 2.9]).
@@
Let us already mention that in the case of $\mathcal{D}=0$, one can derive an equivalent two-step formulation, eliminating $w^{n}$; see Lemma 3.4 below.
@@ 3.2. An implicit–explicit scheme of second order
Note that this system has a unique solution, since the operator in brackets on the left-hand side is elliptic on $\mathcal{V}_{\ker}$ and $\mathcal{B}$ is inf–sup stable; see [AltZ20, Lem 3.1]. Second, we obtain $u^{n+1}\in\mathcal{V}$ by
@@
Since there is no $\mathcal{V}_{\ker}$-elliptic operator in front of $w^{n+1}$ in the first equation, the existence of a unique solution $w^{n+1}\in\mathcal{H}_{\ker}+\mathcal{V}_{\text{c}}$ is only guaranteed if $\mathcal{H}_{\ker}$ is a proper subspace of $\mathcal{H}$; cf. [Zim21, Lem. 9.22]. In this case, the expression $\mathcal{B}w^{n+1}$ is indeed well-defined, even though $w^{n+1}\not\in\mathcal{V}$. In the following, we assume the unique solvability of (3.2d)-(3.2e) such that the IMEX scheme (3.2) is well-defined overall.
In the case where $\mathcal{H}_{\ker}$ is no proper subspace of $\mathcal{H}$, we may replace (3.2d) and (3.2e) by
(3.3) $w^{n+1}=2w^{n+1/2}-w^{n}+\tfrac{\tau}{2}\,\big(f^{n+1}-f^{n}\big)$
and later decompose $w^{n+1}=w^{n+1}_{\ker}+w^{n+1}_{c}$, where we choose $w^{n+1}_{c}=\tfrac{1}{\tau}\,\mathcal{B}^{-}\big(g^{n+3/2}-g^{n+1/2}\big)$. This decomposition, however, is only for theoretical purposes and not needed in the implementation of the method. In addition, we emphasize that the methods are not equivalent in the proper subspace case, but will produce the same approximations $u^{n}$ in the case $\mathcal{D}=0$ and have the same error bounds, as we will explain in detail in Remark 3.10 below.
@@
Remark 3.3.
In the practical implementation of the method, we would first perform a spatial discretization, e.g., by finite elements, leading to matrices $A$, $D$, and $B$ as discrete versions of $\mathcal{A}$, $\mathcal{D}$ and $\mathcal{B}$, respectively. Moreover, we obtain a mass matrix $M$ that appears in front of the second-order terms. Systems (3.2a)–(3.2b) and (3.2d)–(3.2e) then lead to saddle point problems with matrices
$\begin{bmatrix}M+\tfrac{\tau}{2}D+\tfrac{\tau^{2}}{4}A&\tfrac{\tau}{2}B^{T}\\ B\end{bmatrix}\qquad\text{and}\qquad\begin{bmatrix}M&B^{T}\\ B\end{bmatrix}$.
@@
$\quad=-\frac{\tau^{2}}{4}\mathcal{A}\,\big(\tau w^{n+1/2}+\tau w^{n-1/2}\big)-\frac{\tau^{2}}{2}\,\mathcal{A}\big(u^{n}+u^{n-1}\big)+\tau^{2}f^{n}-\tau\mathcal{B}^{*}\big(\tfrac{\tau}{2}\,\lambda^{n+1}+\tfrac{\tau}{2}\,\lambda^{n}+\mu^{n}\big)$
$\quad=-\frac{\tau^{2}}{4}\mathcal{A}\,\big(u^{n+1}+2u^{n}+u^{n-1}\big)+\tau^{2}f^{n}-\tau\mathcal{B}^{*}\big(\tfrac{\tau}{2}\,\lambda^{n+1}+\tfrac{\tau}{2}\,\lambda^{n}+\mu^{n}\big).$
@@ 3.3. Convergence analysis
We introduce $w(t)=\dot{u}(t)$ with its unique decomposition $w=w_{\text{ker}}+w_{\text{c}}$ as before. Note that this implies $w_{\text{c}}(t_{n})=\dot{g}^{n}\coloneqq\dot{g}(t_{n})$. With this, we obtain by (3.2e),
@@
$=w_{\text{ker}}(t_{n})-w_{\text{ker}}^{n}+\mathcal{B}^{-}\bigl(\dot{g}^{n}-\tfrac{1}{\tau}\,(g^{n+1/2}-g^{n-1/2})\bigr)$
@@
We aim to show a similar error representation for the parts in the kernel of $\mathcal{B}$ as in [HocL21]. With $\mathcal{D}_{\ker}$ from Section 2.1, we introduce the operators $\mathcal{R}_{\pm}$ and $\mathcal{R}$ by
These operators appear naturally, when interpreting the IMEX method as perturbation of the Crank–Nicolson method. Here, $\mathcal{R}_{\pm}$ encode the structure of the implicit midpoint rule. However, this formulation is not suited for the implementation, and we refer to [HocL21] for more details. The essential properties of these operators are collected in the following lemma.
Lemma 3.7.
(3.6) $\tau\mathcal{R}_{+}^{-1}\begin{bmatrix}-x\\ \mathcal{D}_{\ker}x\end{bmatrix}=(\operatorname{id}-\mathcal{R})\begin{bmatrix}0\\ x\end{bmatrix}$
and, thus, the bound
Proof.
The bounds on $\mathcal{R},\mathcal{R}_{+}^{-1}$ follow from [HocL21, Lem. 2.4]. From the definition of $\mathcal{R}_{\pm}$, we have
and multiplying with $\mathcal{R}_{+}^{-1}$ gives the claim. ∎
For technical reasons, we introduce the following projection operator.
@@
where we have the two defects $\delta^{n}_{u}$ and $\delta^{n}_{w}$ that essentially come from the trapezoidal rule and satisfy $\|\delta^{n}_{u}\|_{\mathcal{V}_{\ker}}+\|\delta^{n}_{w}\|_{\mathcal{H}_{\ker}}\leq C\tau^{3}$. For $\widehat{f}^{n}=f(t_{n},u(t_{n}))$ and $f^{n}=f(t_{n},u^{n})$, the other terms are given by
@@
We apply $\mathcal{R}_{+}^{-1}$ to (3.8) and resolve the error recursion to
$\begin{bmatrix}e_{u}^{n+1}\\ e_{w}^{n+1}\end{bmatrix}=\mathcal{R}^{n+1}\begin{bmatrix}e_{u}^{0}\\ e_{w}^{0}\end{bmatrix}$ $+\sum\limits_{j=0}^{n}\mathcal{R}^{n-j}\mathcal{R}_{+}^{-1}\begin{bmatrix}0\\ \pi_{\text{ker}}\,S_{2}^{j}\end{bmatrix}+\mathcal{R}^{n-j}\mathcal{R}_{+}^{-1}\begin{bmatrix}\pi_{\text{ker}}\,\delta^{j}_{u}\\ \pi_{\text{ker}}\,\delta^{j}_{w}\end{bmatrix}$
$+\sum\limits_{j=0}^{n}\mathcal{R}^{n-j}\mathcal{R}_{+}^{-1}\begin{bmatrix}\pi_{\text{ker}}\,S_{1}^{j}\\ -\mathcal{D}_{\ker}\pi_{\text{ker}}\,S_{1}^{j}\end{bmatrix}+\mathcal{R}^{n-j}\mathcal{R}_{+}^{-1}\begin{bmatrix}-\pi_{\text{ker}}\,(\delta^{j}_{1}+\delta^{j}_{2})\\ \mathcal{D}_{\ker}\pi_{\text{ker}}\,(\delta^{j}_{1}+\delta^{j}_{2})\end{bmatrix}$.
With the bounds in Lemma 3.7, we observe that the first term accounts for the classical linear stability term, the second term is estimated with the Lipschitz property of $f$, and the third term simply yields a sum over $\mathcal{O}(\tau^{3})$ terms. The fourth and fifth term, on the other hand, need to be taken care of because their upper entries cannot be estimated in $\mathcal{V}_{\ker}$.
By the application of $\mathcal{R}_{+}^{-1}$, however, we can use the relation (3.6) and the term $(\operatorname{id}-\mathcal{R})$ allows for summation by parts. This enables us to replace the spatial regularity in $\mathcal{V}_{\ker}$ by an additional (discrete) time derivative that maintains the order of the defects and the stability term. We refer to the proofs of [HocL21, Th. 2.9 and 3.3], which then also apply to the present case. ∎
@@
We introduce the notation $\widehat{u}_{\text{ker}}^{n}=u_{\text{ker}}(t_{n})$, $\widehat{u}_{\text{c}}^{n}=u_{\text{c}}(t_{n})$ and further denote with $\delta^{n}_{u}$, $\delta^{n}_{w,1}$ the defects from the trapezoidal rule, which satisfy $\|\delta^{n}_{u}\|_{\mathcal{V}}+\|\pi_{\text{ker}}\delta^{n}_{w,1}\|_{\mathcal{H}}\leq C\tau^{3}$. We can thus conclude that the exact solution satisfies
@@
Remark 3.10.
Let us finally comment on the case that $\mathcal{H}=\mathcal{H}_{\ker}$ and the modified method, where $w^{n+1}$ is computed by (3.3). We first note that the same proof as in Lemma 3.4 yields the equivalence of the two schemes if we omit the $\mathcal{B}^{*}\mu$-terms in all calculations. Similarly, in the proof of Lemma 3.9, the computations until (3.9) are fully analogous, again omitting the $\mathcal{B}^{*}\mu$-terms. From that on, we set $w_{\text{c}}^{n+1}=\tfrac{1}{\tau}\,\mathcal{B}^{-}\big(g^{n+3/2}-g^{n+1/2}\big)$ and $w_{\text{ker}}^{n+1}=w^{n+1}-w_{\text{c}}^{n+1}$ (even though $w_{\text{ker}}^{n+1}$ might not be in the kernel of $\mathcal{B}$, but in $\mathcal{H}=\mathcal{H}_{\ker}$). Then, the very same calculations yield the error recursion in (3.8). The estimates of Lemma 3.7 hold on $\mathcal{V}_{\ker}\times\mathcal{H}$ and it is sufficient that $w_{\text{ker}}^{n+1}\in\mathcal{H}$ without any knowledge on the constraint. In particular, we have the same error bound as in Theorem 3.6 under the same assumptions.
@@ 4. A Gautschi-type Integrator
This section is devoted to the extension of Gautschi-type integrators to constrained systems. For this, we restrict ourselves to the damping-free free case, i.e., we set $\mathcal{D}=0$. Moreover, we consider the stronger assumption on the nonlinearity, i.e., Assumption 2.6 (a). This is related to the fact that in the proof of Theorem 4.4 a stronger notation of stability is required to prevent order reduction.
@@ 4.1. Review of the unconstrained case
We recall solution strategies for conservative second-order systems as discussed in [HocL99, Gri02]. Considering the corresponding first-order formulation, we consider the equations
@@
With the solution formula (4.4), we conclude for the solution of (4.1) that
$u(t+\tau)+u(t-\tau)=2\,\cos(\tau\Omega)\,u(t)+2\int_{0}^{\tau}(\tau-s)\,\operatorname{sinc}((\tau-s)\,\Omega)\,f(s,u(s))\,\mathrm{d}s.$
For a constant right-hand side, i.e., for $f(t,u)\equiv f$, this further simplifies to
$u(t+\tau)+u(t-\tau)$ $=2\,\cos(\tau\Omega)\,u(t)+2\,\mathcal{A}^{-1}\big[1-\cos(\tau\Omega)\big]\,f$
@@
Following [HocL99], this may also be written in the alternative form
@@ 4.2. An implicit solution formula
We come back to the constrained case and consider the PDAE (2.1). In the linear case, i.e., if $f(t,u)\equiv f(t)$, we can explicitly write down the solution using the decomposition of $\mathcal{V}$ from (2.3), the first-order formulation, and the variation-of-constants formula; cf. [EmmM13]. This solution formula is also applicable to the general semi-linear case, leading to an implicit formula. As in the unconstrained case, quadrature rules then yield numerical schemes.